A jeweler maker has 20 ounces of a gold alloy worth $160 an ounce and 10 ounce of a silver alloy worth $140 an ounce. How many ounces of another alloy at $124 an ounce must she add to these in order to have a mixture of the three that she can use to make 3-ounce jewelry worth $438 a piece?
Answer by ankor@dixie-net.com(22740)

A jeweler maker has 20 ounces of a gold alloy worth $160 an ounce and 10 ounce of a silver alloy worth $140 an ounce.
How many ounces of another alloy at $124 an ounce must she add to these in order to have a mixture of the three that she can use to make 3-ounce jewelry worth $438 a piece?
:
He doesn't need any of the $124 an ounce to to make this jewelry
:
let x = amt of gold alloy required to make the 3 oz jewelry
then
(3-x) = amt of silver alloy required
:
160x + 140(3-x) = 438
160x + 420 - 140x = 438
160x - 140x = 438 - 420
20x = 18
x = 18/20
x = .9 oz of the $160 per oz alloy
then
3 - .9 = 2.1 oz of the $140 per oz alloy
:
:
Check this
.9(160) + 2.1(140) =
144 + 294 = 438
:
:
Note the jewelry 438/3 = $146 per oz, which is between the $160 and $140 per oz
So no need for $124 per oz stuff